How many pounds of chocolate worth $1.20 a pound must be mixed with 10 pounds of chocolate worth 90 cents a pound to produce a mixture worth $1.00 a pound?
Answer by likaaka(51)

Create a grid to organize the information
Your rows would be $0.90 Chocolate, $1.20 Chocolate, and the Total Chocolate Mix, and your columns would be the Amount of Chocolate in #s and the Total Costs
(This is the best I can do coming up with a grid)
________________amt (#)_________total costs
$0.90_____________10______________9.00
$1.20______________x______________1.20x
Total Mix_________10+x_____________9+1.2x
Using this grid, you see that 10+x=9+1.2x so solve for x
10+x -x=9+1.2x -x
10=9+.2x
10 -9=9+.2x -9
1=.2x
1/.2=(.2x)/.2
x=5, so you need 5lbs of the $1.20 chocolate to produce a mixture worth $1.00